# Fast LiDAR Informed Visual Search in Unseen Indoor Environments

### v8

@@ Abstract
This paper explores the problem of planning for visual search without prior map information. We leverage the pixel-wise environment perception problem where one is given wide Field of View 2D scan data and must perform LiDAR segmentation to contextually label points in the surroundings. These pixel classifications provide an informed prior on which to plan next best viewpoints during visual search tasks. We present LIVES: LiDAR Informed Visual Search, a method aimed at finding objects of interest in unknown indoor environments. A robust map-free classifier is trained from expert data collected using a simple cart platform equipped with a map-based classifier. An autonomous exploration planner takes the contextual data from scans and uses that prior to plan viewpoints more likely to yield detection of the search target. We propose a utility function that accounts for traditional metrics like information gain and path cost and for the contextual information. LIVES is baselined against several existing exploration methods in simulation to verify its performance. It is validated in real-world experiments with single and multiple search objects with a Spot robot in two unseen environments. Videos of experiments, implementation details and open source code can be found at https://sites.google.com/view/lives-2024/home.
@@ I INTRODUCTION
Planning and real-world execution for autonomous visual exploration are receiving significant attention from the robotics community due to their relevance in a wide range of scenarios including inspection [tan2021comprehensive], target search [balta2017integrated, chen2023hybrid], disaster response [delmerico2017active, tomic2012toward, bi2023cure], agricultural monitoring [christiansen2017designing], reconnaissance and surveillance [grocholsky2006cooperative, liang2022reconnaissance, maini2020visibility]. Active planners in unknown spaces are required to make decisions based on incomplete and noisy information about partial environments. This requires planners equipped to generate high quality plans under uncertainty.
@@
Today’s robots are frequently equipped with LiDAR sensors that cast a full view of the surroundings. State of the art exploration planners like [best2022resilient, vidal2020multisensor] demonstrate efficiency by fusing LiDAR and visual information. At the same time, the potential to use additional information from LiDAR scans for online control has grown with advances in parallel computing. While several methods exist for 3D LiDAR scan segmentation [qi2017pointnet, qi2017pointnet++, zamorski2020adversarial], these methods are limited by high computational requirements [jhaldiyal2023semantic]. In contrast, real time projection-based methods often detect moving obstacles to be avoided for autonomous driving applications [chen2022automatic, liu2022less, baur2021slim, mersch2022receding]. Numerous planners are designed to reduce uncertainty in new environments by exploiting features common to places they are designed for [saroya2020online, ramakrishnan2020occupancy].
Figure 1: LIVES allows the Spot robot to search for objects quickly in varied indoor environments. Video of experiments can be found at our website.
Despite improvements in efficiency of planning algorithms for both LiDAR and vision sensors [best2022resilient, dai2020fast, tao2023seer, vidal2020multisensor], it remains an open problem to enhance performance in autonomous planning and execution for the task of visual search [wang2019autonomous]. This work aims to reduce uncertainty in indoor environments by segmenting 2D LiDAR scans online for utilization by a next-best view (NBV) planner.
Figure 2: Overview of a planning step of LIVES. In (a) raw 2D LiDAR scans, pose information and image data are received from the Spot robot. They are sent through a neural network classifier III-C (b) to label each point as map or non-map (c). Separately, LiDAR scans and pose data are used to maintain the global search maps for the two sensors (d). Global search maps are used to generate frontier points, shown in cyan and pink. Frontier centroids are shown as white dots and used for sampling. Each centroid has four viewpoints scored (e) III-E2. In this toy example the left viewpoint in the lower figure in (e) will be scored highest and selected as the next waypoint due to the high number of unknown cells and non-map points observed. Finally, the selected viewpoint is sent back to the robot for execution.
@@
The key insight in this work is to exploit contextual information available in wide Field of View (FoV) LiDAR scans to filter out points unlikely to be the search target. This method deviates from traditional exploration methods, which focus on completeness. Rather, LIVES prioritizes regions more likely to be the search target in the robot’s surroundings. We hypothesize exploring agents can improve time efficiency in visual search by exploiting this additional information. A map based classifier is used to train a model to recognize features inherent to many indoor environments. The perception model feeds the planner the non-permanent points, non-map points, for use during planning. These points guide the NBV planner towards possible search target locations, assumed to be in the set of non-map points in the environment. Priority is given to features of interest using a novel utility function that accounts for both traditional metrics and non-map points. Examples of search environments are shown in Fig. 1. An overview of the proposed method can be found in Fig. 2.
@@
  • A learning strategy for online scan segmentation as map and non-map in unseen, real-world environments
  • A formulation that incorporates classified scan pixel information into a planner for visual search
  • Demonstrate reduced mission completion time in the visual search task over existing planners
@@ II Related Works
II-A Next-Best View and Multisensor Planning
Exploration is often solved using next-best viewpoint planners like Frontier Exploration [yamauchi1997frontier, dai2020fast]. A common approach uses an active perception formulation to maximize entropy reduction [julian2014mutual, charrow2015information]. Ref. [chen2023hybrid] combine frontier and sampling based methods while other methods leverage learning to plan [cao2023ariadne, lodel2022look]. These methods demonstrate efficient exploration in simulation and on hardware, however none are explicitly designed for visual search using contextual information. Ref. [zheng2023system] propose a POMDP for local object search in cluttered 3D environments and extend it with a 2D planner for larger areas, however, their work is aimed at 3D search for small objects, rather than quickly searching large areas.
Multisensor exploration includes LiDAR information in planning for visual exploration tasks. State of the art multisensor exploration with multiple UAVs [best2022resilient] leverages the single agent exploration algorithm presented in [vidal2020multisensor]. These methods combine frontier generation for both sensors and use a single utility function to score frontier candidates for surface mapping. Ref. [patel2022lidar] maintains a dense 3D pointcloud for recognizing of objects in subterranean tunnel environments, which are used by a Pan-Tilt-Zoom camera. These works demonstrate efficiency in a number of real world environments, however none of these methods is designed for the environments considered in this article.
II-B World Prediction Based Planning
World prediction methods typically learn a model of relevant environments and exploit their structure to explore high-value regions first. Ref. [saroya2020online] use topological features from a database of tunnel networks to inform a frontier-based exploration policy. Similarly, ref. [tao2023seer] use a prediction module to detect semantic objects and classify frontiers, resulting in exploring rooms first before continuing global exploration. The proposed method instead aims to ignore low value regions in order to reduce search completion time. Prediction guide the robot towards likely exit locations.
II-C LiDAR Classification
Indoor 3D pointcloud classification is tackled in [thomas2021self] with KPConv [thomas2019kpconv]. This requires the offline annotation and map building at instantiation and is used for simulated localization and navigation [thomas2022learning]. The proposed work instead focus on low computation 2D scan pixel classification.Ref. [pomerleau2013comparing] presents an in depth comparison and framework for Iterative Closest Point (ICP) methods, similar to that of [thomas2021self], focused on map building. Finally, [shan2018lego] present a map-free framework for segmenting 3D pointclouds into ground, edges and planar pointsfor mapping outdoor environments.
@@ III Methods
The overall scan classifier and planning module are described in Fig. 2 and Alg. 1. The key components are detailed in this section.
@@ III-A Environment
The environment is described as a discrete set of points $\mathcal{E}$. This can be split into two subsets $\mathcal{E}^{\text{free}}\subset\mathcal{E}$ and $\mathcal{E}^{\text{occ}}\subset\mathcal{E}$, which represent the free and occupied cells in the environment. Both of these sets are initially unknown to the robot, but are assumed to follow a structure inherent in indoor environments (e.g. hallways, planar walls, open rooms and loops).
@@ III-B Ground Truth LiDAR Scan Classification
The search target of interest belongs to $\mathcal{E}^{\text{occ}}\subset\mathcal{E}$. In particular, scan classification attempts to divide $\mathcal{E}^{\text{occ}}$ into two subsets, namely $\mathcal{E}^{\text{non-map}}$ and $\mathcal{E}^{\text{map}}$. In this work, $\mathcal{E}^{\text{map}}$ are points in the environment that are deemed to be Long-Term Features (LTFs) during classification, while $\mathcal{E}^{\text{non-map}}$ are points that are either Short-Term Features (STFs) or Dynamic Features (DFs) [biswas2017episodic]. LTFs represent the static map obstacles, STFs represent static un-mapped obstacles and DFs represent moving points. Given map information, scan points can be classified into these three categories as follows.
Let $x_{i}$ denote the pose of the robot, and $s_{i}$ denote observation at time step $t_{i}$. Each observation $s_{i}$ consists of $n_{i}$ 2D points, $s_{i}=\{p^{j}_{i}\}_{j=1:n_{i}}$. Observations are transformed from robot local frame into the global frame using an affine transformation $T_{i}\in SE(3)$. Finally, let map $\mathcal{M}$ be represented as a set of lines $\{l_{i}\}_{1:n}$.
@@ III-B1 LTF
First, an analytic ray cast is performed [biswas2012depth] to determine expected laserscan based on map $\mathcal{M}$ and current robot position $x_{i}$. Given observations, the probability that points correspond to one of the lines of that static map can be written
@@
where $\Sigma_{s}$ is the scalar variance of observations, which comes from sensor accuracy. If Eq. 1 is greater than a threshold, point $p_{i}^{j}$ is classified as a LTF.
@@ III-B2 STF
Remaining points will be classified as STF or DF. Observations at current time $i$, $p_{i}^{j}$, are compared with prior observations at time $k$, $p_{k}^{l}$ to determine correspondence between points in subsequent observations. The likelihood of the remaining points corresponding to the same point as in a previous laserscan is computed as
@@
When Eq. 2 is greater than some threshold, point $p_{i}^{j}$ is classified as an STF. Remaining points in $p_{i}^{j}$ are classified as DFs. The result is classified scan $s^{\text{class}}_{i}$.
@@ III-C Map-free LiDAR Scan Classification
The classification method defined in III-B is limited by the necessity of static map information $\mathcal{M}$. Therefore, this work uses supervised learning to train a model to reproduce classifications for $\{p^{j}_{i}\}_{j=1:n_{i}}$ in scan $s_{i}$.
@@ III-C1 Dataset and Data Acquisition
The dataset $\mathcal{D}$ consists of tuples of data from the ground truth classification model. The tuples are given as $\mathcal{D}=\{(x_{i},s_{i},s_{i}^{\text{class}})\}^{N}_{i=1}$, where $x_{i}$ represents robot pose, $s_{i}$ is the raw LiDAR scan data, and $s_{i}^{\text{class}}$ is the classified scan. $N$ is the total number of data points collected. Data acquisition with a robot requires a human expert to control, which is difficult to scale due to mechanical limits of the robot. To ease data acquisition we design a steerable cart platform (see website for details) to mimic the robot. Human operators easily maneuver this cart to obtain the dataset. We mount a Velodyne Puck LiDAR for recording scan information and a front-mounted Intel RealSense T265 stereo camera to obtain the cart’s odometry estimate, required for running the ground-truth classification. A laptop maintains the pose estimate [biswas2017episodic] and collects the data. The Puck has $n_{i}=897$ and classified scans are recorded at 5Hz. A dataset of size $N\simeq 145,000$ is collected in this study, representing roughly 8 hours of time. In practice, the model only receives raw scan data, pose estimate and previously predicted labels from the robot during operation and must classify pixel-wise the scan online. Training data can be found on the project website.
@@ III-C2 Architecture and Training
We formulate pixel-wise LiDAR classification as a supervised learning problem during training and an auto-regressive problem during inference. Given dataset $\mathcal{D}$, we shuffle and batch the dataset between time steps and locations to minimize location-based and temporal bias during training. Because no map information is provided explicitly, we provide the model with a history buffer of $k$ time steps for scan ranges and robot poses, and $k-1$ previously estimated ground truth labels. Because we do not have access to ground truth labels at inference time, we corrupt the ground truth labels during training by randomly bit-flipping the classification of 10% of the labels, chosen by sampling indices from a uniform random distribution. We then take the exponential weighted average of the last $k-1$ corrupted pixel-wise classifications to estimate the current pixel-wise classification. This estimate is concatenated with the $k-1$ corrupted ground truth labels to produce a $k$-length matrix of estimated scan labels. During inference, the label history buffer is populated by previous estimates from the policy. The history buffer is concatenated with the updated pixel-wise exponential weighted average. At initialization, the policy is bootstrapped with zero-value poses, scans, and labels, and run for $k$ steps of inference until the history buffer is full. Code for training a model can be found on the project website.
@@
Figure 3: Pixel-wise LiDAR scan classification model architecture. This model enables speeding up search by providing contextual information to the planner. $k$ is the length of the history buffer. (a) $[3,k]$ pose history matrix containing $[x,y,\theta]$. (b) $[n_{i},k]$ LiDAR range history matrix. (c) $[n_{i},k-1]\cup[n_{i},1]$ estimated label history matrix concatenated with its pixel-wise exponential weighted average. (d) The model consists of three temporal-convolutional encoders (TCN), one for each input. The encoded poses, scans, and labels are combined to produce a pixel-wise classification of the LiDAR scan. In (e) a threshold (positive/negative) is applied to the raw logits such that each pixel is classified as either a map point or a non-map point.
@@
The full model architecture is shown in Fig. 3. The policy consists of a temporal-convolutional encoder (TCN) for each input. Each encoder contains a single convolutional layer with scan-wise circular padding and a single linear layer. The scan and label encoders have kernel size $[k,k]$, and the pose encoder has kernel size $[1,3]$. Hyperbolic tangent activations are used for all layers. The output of each encoder is the same size as the LiDAR scan to be classified. The output of the pose encoder is summed with the output of the scan encoder as a pose correction, and then normalized by applying a hyperbolic tangent function. The resulting LiDAR scan encoding is multiplied element-wise with the label encoding. A threshold (positive/negative) is applied to the raw logits such that each pixel is classified as either a map point or a non-map point. The policy is trained with mean squared error loss against ground truth classifications. Training for 20 epochs on a workstation Nvidia 3080 12Gb takes 3 minutes.

### v11

@@ Abstract
This paper details a system for fast visual exploration and search without prior map information. We leverage frontier based planning with both LiDAR and visual sensing and augment it with a perception module that contextually labels points in the surroundings from wide Field of View 2D LiDAR scans. The goal of the perception module is to recognize surrounding points more likely to be the search target in order to provide an informed prior on which to plan next best viewpoints. The robust map-free scan classifier used to label pixels in the robot’s surroundings is trained from expert data collected using a simple cart platform equipped with a map-based classifier. We propose a novel utility function that accounts for the contextual data found from the classifier. The resulting viewpoints encourage the robot to explore points unlikely to be permanent in the environment, leading the robot to locate objects of interest faster than several existing baseline algorithms. Our proposed system is further validated in real-world search experiments for single and multiple search objects with a Spot robot in two unseen environments. Videos of experiments, implementation details and open source code can be found at https://sites.google.com/view/lives-2024/home.
@@ I INTRODUCTION
Planning and real-world execution for autonomous visual exploration are receiving significant attention from the robotics community due to their relevance in a wide range of scenarios including inspection [nguyen2024grey], search [chen2023hybrid], disaster response [bi2023cure], agricultural monitoring [christiansen2017designing], reconnaissance and surveillance [liang2022reconnaissance]. As mobile robots begin to proliferate in our communities, visual planning capabilities remain fundamental for autonomous agents to achieve success in human environments [zheng2023system]. Active planners in unknown spaces are required to make decisions based on incomplete and noisy information about partial environments. This requires planners equipped to generate high quality plans under uncertainty.
@@
Today’s robots are frequently equipped with LiDAR sensors that cast a full view of the surroundings. State of the art visual exploration planners like [best2022resilient, vidal2020multisensor] demonstrate efficiency by fusing LiDAR and visual information. At the same time, the potential to detect additional information in LiDAR scans for online control has grown with advances in parallel computing. While methods exist for 3D LiDAR segmentation [qi2017pointnet, zamorski2020adversarial], they are limited by high computational requirements [jhaldiyal2023semantic]. In contrast, projection-based methods run online but are used to detect moving obstacles for autonomous vehicles [chen2022automatic, mersch2022receding]. The literature also demonstrates planning under reduced uncertainty in unseen environments by exploiting features common to the places they are designed for [saroya2020online, ramakrishnan2020occupancy]. Despite improvements in efficiency of planning algorithms for both LiDAR and vision sensors [best2022resilient], it remains an open problem to enhance performance in autonomous planning and execution [wang2019autonomous]. To this end, we propose to exploit features common to many indoor environments using a lightweight scan segmentation method for efficient visual search planning.
@@
The key insight in this work is to exploit contextual information available in wide Field of View LiDAR scans to filter out points unlikely to be the search target. This deviates from traditional exploration methods, which focus on completeness. Instead, the proposed method prioritizes regions more likely to be the search target in the robot’s surroundings. We hypothesize exploring agents can improve time efficiency in visual search by exploiting this additional information. A map based classifier is used to train a model to recognize features inherent to many indoor environments. The perception model feeds the planner the non-permanent points, non-map points. These points guide the Next Best View (NBV) planner towards possible search target locations, assumed to be in the set of non-map points in the environment. Examples of search environments are shown in Fig. LABEL:fig:spot.
@@
  • Achieve significantly faster search indoors than baselines by augmenting multisensor exploration with contextual information from LiDAR scans
  • A scan segmentation policy that identifies non-permanent points in unseen, real-world environments
  • A novel frontier-based exploration formulation that accounts for segmented scan information
@@ II Related Works
The two common approaches to exploration include frontier [yamauchi1997frontier] and information-theoretic sampling methods [charrow2015information]. Sampling approaches avoid the expensive frontier generation process but require evaluation of every sample, motivating hybrid approaches [dai2020fast]. Alternatively, [umari2017autonomous] use RRTs to search for frontier points; achieving efficiency using multiple trees and decoupling tree expansion from movement. Another approach expands RRTs with accessible space and executes the edge with the highest information gain in a receding horizon fashion [bircher2016receding]. The key difference between these methods and ours is the addition of segmented scans, which is shown to significantly improve task completion time.
State of the art multisensor exploration is shown with an underwater vehicle [vidal2020multisensor] and multiple UAVs [best2022resilient]. They combine frontier generation for two sensors and use a single utility function to score frontier candidates for surface inspection. We modify their method to sample candidates at frontier centroids instead of surrounding them. This modified method is equivalent to our proposed planner without the use of segmented scan information and represents a third baseline method. This baseline clearly demonstrates the positive effect of the proposed contextual information.
Partially observable markov decision processes (POMDP) are commonly used in visual object search [kaelbling1998planning, zheng2023system]. The latter presents a platform-generalizable 3D viewpoint planner, capable of reasoning about viewpoints underneath furniture or between appliances. Their work includes a 2D planner for multi-room search, however, it is not tailored to search larger spaces for things that may be in plain sight; a skill that offers real-world application in search and rescue. Because our goals involve searching larger spaces quickly, the aforementioned exploration algorithms provide more fair baseline comparisons for the proposed system.
World prediction is used to learn a model of relevant environments and exploit their structure to explore high-value regions first. Ref. [saroya2020online] use topological features like loops or dead-ends from a database of subterranean tunnels to inform a frontier based exploration policy. Instead, [ramakrishnan2020occupancy] considers occupancy anticipation from egocentric RGB-D for navigation by an exploration planner. Ref. [caley2019deep] learns to predict locations of exits in building with a convolutional neural network using a database of building blueprints. Our proposed system leverages a similar principle to exploit the structures inherent to many indoor environments - like planar walls, hallways, rooms and loops - in order to guide the planner. We achieve this in a novel way by segmenting scans to identify non-permanent features in the environment. Ultimately, the addition of these features are shown to enable our method to find objects of interest more quickly than the examined baselines.
@@ III Methods
The overall scan classifier and planning module are described in Alg. 1 with key components detailed in this section.
@@ III-A Environment
The environment is a discrete set of points $\mathcal{E}$, split into two subsets $\mathcal{E}^{\text{free}}\subset\mathcal{E}$ and $\mathcal{E}^{\text{occ}}\subset\mathcal{E}$, representing free and occupied cells. Both sets are initially unknown to the robot, but are assumed to follow a structure inherent to indoor environments (e.g. hallways, planar walls, open rooms and loops).
@@ III-B Ground Truth LiDAR Scan Classification
Ground truth classification uses map $\mathcal{M}$ to divide $\mathcal{E}^{\text{occ}}$ into $\mathcal{E}^{\text{non-map}}$ and $\mathcal{E}^{\text{map}}$. In this work, $\mathcal{E}^{\text{map}}$ are points in the environment that are deemed to be Long-Term Features (LTFs) during classification, while $\mathcal{E}^{\text{non-map}}$ are points that are either Short-Term Features (STFs) or Dynamic Features (DFs) [biswas2017episodic]. LTFs represent permanent features while STFs and DFs are non-permanent static and dynamic points, respectively. Given $\mathcal{M}$, represented as a set of lines $\{l_{i}\}_{1:n}$, points are classified into these categories as follows.
Let $x_{i}$ denote robot pose, and $s_{i}$ denote observation at time step $t_{i}$. Observation $s_{i}$ consists of $n_{i}$ 2D points, $s_{i}=\{p^{j}_{i}\}_{j=1:n_{i}}$. Observations are transformed from local to global frame using affine transformation $T_{i}\in SE(3)$.
@@ III-B1 LTF
First, an analytic ray cast is performed [biswas2012depth] to determine expected laserscan based on map $\mathcal{M}$ and current robot position $x_{i}$. Given observations, the probability that points correspond to the static map can be written:
@@
where $\Sigma_{s}$ is the scalar variance of observations from sensor accuracy. If Eq. 1 is above a threshold, point $p_{i}^{j}$ is an LTF.
@@ III-B2 STF
Remaining points are either STFs or DFs. Observations at current time $i$, $p_{i}^{j}$, are compared with prior observations at time $k$, $p_{k}^{l}$ to determine correspondence between points in subsequent observations. The likelihood of a point corresponding between time steps is computed as:
@@
When Eq. 2 is greater than some threshold, point $p_{i}^{j}$ is classified as an STF. Remaining points in $p_{i}^{j}$ are classified as DFs. The result is classified scan $s^{\text{class}}_{i}$. Open source code is published by Biswas et. al. [biswas2017episodic] at [enml].
@@ III-C Map-free LiDAR Scan Classification
The classification method defined in III-B is limited by the necessity of static map information $\mathcal{M}$. Therefore, this work uses supervised learning to train a model to reproduce binary classifications for $\{p^{j}_{i}\}_{j=1:n_{i}}$ in scan $s_{i}$, where $\{p^{j}_{i}\}_{j=1:n_{i},class}\in\{-1,1\}$
@@ III-C1 Dataset and Data Acquisition
We first collect a dataset, however data acquisition with a robot is difficult to scale due to its mechanical limits. To ease data acquisition we design an easily maneuverable cart platform (see fig. 1) to mimic the robot. Data is labeled and recorded online as the cart is pushed through eight indoor environments on the university campus that vary in size and shape in attempt to capture generic features found in different indoor environments. Training and test maps can be found on the website. A laptop with static map information on board the cart receives odometry and scan data from sensors, then uses the method described in the previous section to maintain a pose estimate and collect dataset $\mathcal{D}$, consisting of tuples of data. The tuples are given as $\mathcal{D}=\{(x_{i},s_{i},s_{i}^{\text{class}})\}^{N}_{i=1}$, where $x_{i}$ represents robot pose, $s_{i}$ is the raw LiDAR scan data, and $s_{i}^{\text{class}}$ is the classified scan. Note that non-map elements in $s_{i}^{\text{class}}$ are value $-1$ and map elements are value $1$. $N$ is the total number of data points collected.
Figure 1: The cart used for labeled data acquisition and the Spot robot used for deployment. The RealSense provides odometry estimate to the cart, required for ground-truth estimation. Spot is equipped with an RGB-D Azure Kinect for detection.
The Velodyne Puck has $n_{i}=897$ and classified scans are recorded at 5Hz. A dataset of size $N\simeq 145,000$ is collected in this study, representing roughly 8 hours of time. Training data can be found on the project website.
@@ III-C2 Architecture and Training
In practice, the model only receives raw scan data, pose estimates and previously predicted labels from the robot during operation and must classify pixel-wise the scan online. As a result, we formulate pixel-wise classification as a supervised learning problem during training and an auto-regressive problem during inference. Given dataset $\mathcal{D}$, we shuffle and batch the dataset between time steps and locations to minimize location-based and temporal bias during training. Because no map information is provided explicitly, we provide the model with a history buffer of $k$ time steps for scan ranges and robot poses, and $k-1$ previously estimated ground truth labels. We corrupt the ground truth labels during training by randomly bit-flipping the classification of 10% of the labels, chosen by sampling indices from a uniform random distribution. In order to produce a $k$ length estimated label history input, we take the exponential weighted average of the $k-1$ corrupted pixel-wise classifications to estimate the current timestep’s classification.
During inference, the label history buffer is populated by previous estimates from the policy, cropped by a length $k-1$ sliding-window. The history buffer is concatenated with the updated pixel-wise exponential weighted average in the same way as was done during training to obtain the $k$ length estimated label history input. At initialization, the policy is bootstrapped with zero-value poses, scans, and labels, and run for $k$ steps of inference until the history buffer is full. Code for training a model can be found on the project website.
@@
Figure 2: Pixel-wise LiDAR scan classification model architecture used to speed up search by providing information to the planner. $k$ is the length of the history buffer. (a) $[3,k]$ pose history matrix containing $[x,y,\theta]$. (b) $[n_{i},k]$ LiDAR range history matrix. (c) $[n_{i},k-1]\cup[n_{i},1]$ estimated label history matrix concatenated with its pixel-wise exponential weighted average. (d) The model consists of three temporal-convolutional encoders (TCN) (i.e. pose encoder, scan encoder, and label encoder). The encoded poses, scans, and labels are combined to produce a pixel-wise classification of the LiDAR scan. In (e) a threshold (positive/negative) is applied to the raw logits such that each pixel is classified as either a map point or a non-map point.
@@
The full model architecture is shown in Fig. 2 and consists of a temporal-convolutional encoder (TCN) for each input. Each TCN contains a single convolutional layer with scan-wise circular padding and a single linear layer. The scan and label encoders have kernel size $[k,k]$, and the pose encoder has kernel size $[1,3]$. Hyperbolic tangent activations are used for all layers. The output of each TCN is the same size as the LiDAR scan to be classified ($n_{i}=897$). The output of the pose history TCN is summed with the output of the scan encoder as a pose correction before element-wise multiplication with the label encoding, which may flip the sign of the term. The result is normalized by applying a hyperbolic tangent function such that each value is $\{-1,1\}$. A threshold (positive/negative) is applied to the raw logits such that each pixel is classified as either a map point or a non-map point. The policy is trained with mean squared error loss against ground truth classifications. Training for 20 epochs on a workstation Nvidia 3080 12Gb takes 3 minutes.
Figure 3: Map updates are performed from observation at each timestep for visual (a) and LiDAR (b) sensors. White cells are known free and blue/grey cells are unknown. Next, frontiers and their centroids are computed for each of the two sensors (c) and (d).